OpenStudy (anonymous):

Simplify the rational expression. 1. x+3/x^2-2x-15= 2. x-6/x^2-7x+6= 3. x-2/x^2-8x+12= 4. x+4/x^2+6x+8=

OpenStudy (anonymous):

1.) x² - 2x -15 = 0 x1,2 = + 1 +- sqrt ( 1 +15) x1,2 = 1 +- 4 ... x1 = -3 , x2 = 5 (x+3) * (x -5) x+3 / (x+3 ) * (x - 5) = 1 / (x -5) ...and you have to do this all the time for each instance